# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

### v3

@@ 5.3 Potential for VLA Model Pretraining
A common paradigm in Vision-Language-Action (VLA) models is to initialize the visual backbone using encoders pre-trained on large-scale datasets (e.g., CLIP, VAEs). In this experiment, we investigate whether fine-tuning a pre-trained visual backbone on robot-oriented video data can enhance its suitability for robotic policy learning. Specifically, we fine-tune the pre-trained VAE encoder and decoder used in the Unified Video Action Model(UVA) [li2025unifiedvideoactionmodel] separately on both the original SSv2 and the H2R-augmented datasets. The fine-tuned VAEs are subsequently frozen and integrated into the UVA framework for policy training on two distinct dual-arm UR5 robot tasks(shown in Figure  5). The results shown in Table  11 demonstrate a clear performance trend in the different variants of the UVA model.
Table 11: Success rates of UVA variants with different visual backbones.
Model PlaceBowl StackBlocks Average UVA 0.2 0.2 0.2 UVA + SSv2 0.1 0.2 0.15 UVA + SSv2 + H2R 0.4 0.3 0.35
5.4 Ablation Study
@@
To evaluate the effectiveness of each component in H2R, we conduct ablation studies on two time-consuming steps: (1) performing hand inpainting without overlaying a robotic arm (H2R w/o Overlay), and (2) overlaying the arm without precise alignment between the hand and the camera, instead using random pasting (H2R w/o Retarget). Table 9 shows the necessity and effectiveness of each component in H2R. The first step leads to a significant drop in success rate due to the loss of critical human-object interaction pixels after inpainting. The second step fails to provide accurate motion cues for the model and introduces visual mismatches with real-world manipulation tasks.
6 Conclusion
@@
[…]
Supplementary Material
Appendix A Details of Simulator Camera Position Alignment
We define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robot arm in the simulator. We build the coordinate system $^{W}\mathbf{I}_{H}$ based on the hand keypoints:
$^{W}\mathbf{I}_{H}=\{^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}\}$ (2)
Where $^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}$ are unit vectors along the x-axes, y-axes and z-axes of the human hand coorinate system. With the keypoints get in HaMeR, we build the three axis of coordinates with the following functions:
${}^{w}\mathbf{i}_{H,x}$ $=^{w}\mathbf{i}_{0,9}$ (3)
${}^{w}\mathbf{i}_{H,y}$ $=^{w}\mathbf{i}_{0,9}\times^{w}\mathbf{i}_{0,13}$
${}^{w}\mathbf{i}_{H,z}$ $=^{w}\mathbf{i}_{H,x}\times^{w}\mathbf{i}_{H,y}$
Where $^{w}\mathbf{i}_{0,9}$ and $^{w}\mathbf{i}_{0,13}$ are unit vectors along the middle and ring fingers, respectively. In this notation, the first index (0) refers to the specific finger (middle or ring), and the second index (9 and 13) corresponds to the joint numbers along those fingers, as defined by the MANO model. Similarly, To construct the mapping from hand pose to robot arms, we need to get another coordinate system $^{W}\mathbf{I}_{S}$ in the simulator:
$^{W}\mathbf{I}_{S}=\{^{w}\mathbf{i}_{S,x},^{w}\mathbf{i}_{S,y},^{w}\mathbf{i}_{S,z}\}$ (4)
The method of determining the axis of coordinates is the same:
${}^{w}\mathbf{i}_{S,x}$ $=^{w}\mathbf{i}_{0,2}$ (5)
${}^{w}\mathbf{i}_{S,y}$ $=^{w}\mathbf{i}_{0,2}\times^{w}\mathbf{i}_{0,3}$
${}^{w}\mathbf{i}_{s,z}$ $=^{w}\mathbf{i}_{S,x}\times^{w}\mathbf{i}_{S,y}$
Where $\mathbf{i}_{0,2},\mathbf{i}_{0,3}$ are unit vectors along robot fingers that correspond to human middle and ring fingers and the index corresponds to the joint numbers defined by MANO. We build the following two coordinate transformation matrix to construct the mapping:
${}^{W}_{H}\mathbf{R}$ $=\begin{pmatrix}{}^{W}\mathbf{I}_{H}&\mathbf{key}_{human}\\ \mathbf{O}&1\end{pmatrix}$ (6)
${}^{W}_{S}\mathbf{R}$ $=\begin{pmatrix}{{}^{W}\mathbf{I}_{S}}&{\mathbf{key}_{robot}}\\ \mathbf{O}&{1}\end{pmatrix}$
Where $\mathbf{key}_{human},\mathbf{key_{robot}}$ are the positions of human wrist and robot wrist. After obtaining the two coordinate systems, we need to determine the position of the camera in the simulator ($^{W}\mathbf{cam}_{sim}$) and the position of the camera in the real world ($^{H}\mathbf{cam}_{Real}$), thus we can ensure we get the same pose of the human hand and robot arms
${}^{H}\mathbf{cam}_{Real}$ $=^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (7)
${}^{S}\mathbf{cam}_{sim}$ $=^{H}\mathbf{cam}_{Real}$
${}^{W}\mathbf{cam}_{sim}$ $=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$
Appendix B Details of Experimental Setup
Figure 6: Simulation benchmark. We choose 3 tasks from the Robomimic, 3 tasks from the RLBench, and 3 tasks from the CortexBench, covering a range of robotic manipulation skills. We also include the PushT task, designed for the Diffusion Policy framework, as an additional benchmark to evaluate performance in a different task setup.
@@ Appendix B Details of Experimental Setup
Data Collection in Real-world Experiment. For the UR5-gripper, demonstration trajectories were collected via keyboard-based teleoperation. For the UR5-Leaphand, demonstrations were gathered using vision-guided teleoperation. For the Dual-arm Franka and Dual-arm UR5 setups, data collection was performed through homogeneous-arm bilateral teleoperation.
Additional Evaluation Details. Figure  6 provides visualizations of the different simulation benchmarks, and Figure  7 illustrates the real-world experimental setup. For the simulation benchmarks, we use the default configurations for both training and evaluation. In the real-world tasks, the UR5–Gripper and UR5–Leaphand setups use the camera shown in Figure  7 as the visual input, whereas the dual-arm Franka and dual-arm UR5 configurations use a front–top camera together with two wrist-mounted cameras (three cameras in total) for visual observation. To ensure that the initial conditions are evaluated consistently across algorithms, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstration collection. During evaluation, the object placement region is kept identical to that used in data collection, ensuring a fair and consistent comparison.
@@
Appendix C Failure Case Analysis
Table 12: Sub-goal.Task-specific sub-goal evaluation. To gain fine-grained insights into policy performance, we design a manual rubric covering key sub-goals for each manipulation task. Each cell reports the number of successful vs. unsuccessful attempts (Y/N) over 20 evaluation trials. Results show that models enhanced with H2R consistently accomplish more sub-goals across tasks compared to their baseline counterparts, demonstrating improved robustness in real-world execution. Bold numbers indicate better performance between paired models.
@@ Appendix C Failure Case Analysis
To enable fine-grained evaluation of policy performance and gain deeper insights into failure cases, we designed a task-specific evaluation rubric. Table 12 displays our rubric that the evaluator filled out when rolling out different policies. Take the DP policy as an example, the results in Table 12 demonstrate that H2R-augmented visual representation models not only improve overall success rates in real-world tasks, but also allow to accomplish more than half of the task consistently.
Table 13: Robomimic Experiment result. We report the success rate (%) over IL-based tasks for MAE and R3M Robomimic.
Table 14: Generalization under Lighting Variations. Success rates (%) under real-world lighting disturbances. Models trained with additional lighting augmentations (LightAug) show significant improvements in robustness compared to baseline and H2R-only models.
Tasks MAE MAE+H2R MAE+H2R+LightAug R3M R3M+H2R R3M+H2R+LightAug Leaphand-GraspChicken 10 10 20 0 0 10 Leaphand-StandCup 20 25 40 5 15 20 Leaphand-OpenBox 5 0 25 10 10 10 Average 11.7 11.7 28.3 5 8.3 13.3
Appendix D Additional Ablation Study in Simulation Experiment
@@
From Table 13, we observe that the encoder trained on H2R processed data shows consistent improvements across various tasks compared to the encoder trained on the original data, with the average success rate on all tasks ranging from 0.9% to 10.2%. Especially for the more challenging MoveCan task, it can improve the success rate by 25.5%. Additionally, while encoders trained on the relatively simple CutMix data show improvement on tasks in Robomimic, their performance in the PushT task remains slightly worse than the encoders trained on original data. These results demonstrate the effectiveness of using the robotic arm to cover the human hand in video data, as well as the effectiveness of H2R in imitation learning.
Appendix E Generalization on Light Condition.
Figure 9: Lighting Setup in Real-World Experiments. Light configuration used in real-world evaluations.
To evaluate the generalization under varying lighting conditions, we introduce illumination disturbances during evaluation, as illustrated in Figure 9. Additionally, during training, we incorporate randomized lighting with varying directions and colors into the simulation environment for data augmentation. We compare three settings: no augmentation(MAE, R3M), H2R augmentation(MAE+H2R, R3M+H2R), and H2R with lighting disturbances(MAE+H2R+LightAug, R3M+H2R+LightAug). As shown in Table 14, the model trained with H2R and lighting perturbations demonstrates significantly better generalization to real-world lighting variations than other baselines, highlighting the effectiveness of H2R in bridging the domain gap caused by lighting variations.

### v4

@@ IV-C Real-world Results
Synergy with Vision-Language-Action Models. A common paradigm in Vision-Language-Action (VLA) models is to initialize the visual backbone using encoders pre-trained on large-scale datasets (e.g., CLIP, VAEs). In this experiment, we investigate whether fine-tuning a pre-trained visual backbone on robot-oriented video data can enhance its suitability for robotic policy learning. Specifically, we fine-tune the pre-trained VAE encoder and decoder used in the Unified Video Action Model(UVA) [32] separately on both the original SSv2 and the H2R-augmented datasets. The fine-tuned VAEs are subsequently frozen and integrated into the UVA framework for policy training on two distinct dual-arm UR5e robot tasks. The results shown in Table  X demonstrate a clear performance trend in the different variants of the UVA model.
TABLE X: UVA with different visual backbones. Success rates ($\uparrow$) on dual-arm UR5e tasks when using UVA with original backbone, and with VAEs fine-tuned on SSv2 or H2R-augmented data.
Method PlaceBowl StackBlocks Avg. UVA 0.20 0.20 0.20 UVA (SSv2) 0.10 0.20 0.15 UVA (H2R) 0.40 0.30 0.35
TABLE XI: Ablation Study. Ablation results by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
Policy Method GraspChicken StandCup OpenBox Avg. DP H2R 55.0 60.0 65.0 60.0 H2R w/o Overlay 30.0 40.0 20.0 30.0 Gain ($\Delta$) -25.0 -20.0 -45.0 -30.0 H2R w/o Retarget 30.0 55.0 45.0 43.3 Gain ($\Delta$) -25.0 -5.0 -20.0 -16.7 ACT H2R 50.0 50.0 40.0 46.7 H2R w/o Overlay 25.0 35.0 25.0 28.3 Gain ($\Delta$) -25.0 -15.0 -15.0 -18.3 H2R w/o Retarget 45.0 30.0 15.0 30.0 Gain ($\Delta$) -5.0 -20.0 -25.0 -16.7
@@
Ablation Study. To evaluate the effectiveness of each component in H2R, we conduct ablation studies on two time-consuming steps: (1) performing hand inpainting without overlaying a robotic arm (H2R w/o Overlay), and (2) overlaying the arm without precise alignment between the hand and the camera, instead using random pasting (H2R w/o Retarget). Table XI shows the necessity and effectiveness of each component in H2R. The first step leads to a significant drop in success rate due to the loss of critical human-object interaction pixels after inpainting. The second step fails to provide accurate motion cues for the model and introduces visual mismatches with real-world manipulation tasks.
TABLE XII: Generalization under Lighting Variations. Success rates ($\uparrow$) on UR5-Leaphand tasks under real-world lighting disturbances, comparing no augmentation, H2R, and H2R with lighting augmentation (H2R+LightAug).
Method GraspChicken StandCup OpenBox Avg. MAE 10.0 20.0 5.0 11.7 MAE (H2R) 10.0 25.0 0.0 11.7 Gain ($\Delta$) 0.0 +5.0 -5.0 0.0 MAE (H2R + LightAug) 20.0 40.0 25.0 28.3 Gain ($\Delta$) +10.0 +20.0 +20.0 +16.6 R3M 0.0 5.0 10.0 5.0 R3M (H2R) 0.0 15.0 10.0 8.3 Gain ($\Delta$) 0.0 +10.0 0.0 +3.3 R3M (H2R + LightAug) 10.0 20.0 10.0 13.3 Gain ($\Delta$) +10.0 +15.0 0.0 +8.3
Robustness to Lighting Variations. To evaluate the generalization under varying lighting conditions, we introduce illumination disturbances during evaluation. Additionally, during training, we incorporate randomized lighting with varying directions and colors into the simulation environment for data augmentation. We compare three settings: no augmentation(MAE, R3M), H2R augmentation(H2R), and H2R with lighting disturbances(H2R+LightAug). As shown in Table XII, the model trained with H2R and lighting perturbations demonstrates significantly better generalization to real-world lighting variations than other baselines, highlighting the effectiveness of H2R in bridging the domain gap caused by lighting variations.
V Conclusion
@@
Appendix A Details of Experimental Setup
@@
TABLE XIII: Robomimic Experiment result. We report the success rate (%) over IL-based tasks for MAE and R3M Robomimic.
A-A Additional Evaluation Details.
Figure  3 provides visualizations of the different simulation benchmarks, and Figure  4 illustrates the real-world experimental setup.
@@
Appendix B Failure Case Analysis
TABLE XIV: Task-specific sub-goal evaluation. To gain fine-grained insights into policy performance, we design a manual rubric covering key sub-goals for each manipulation task. Each cell reports the number of successful vs. unsuccessful attempts (Y/N) over 20 evaluation trials. Results show that models enhanced with H2R consistently accomplish more sub-goals across tasks compared to their baseline counterparts, demonstrating improved robustness in real-world execution. Bold numbers indicate better performance between paired models.
@@ Appendix B Failure Case Analysis
To enable fine-grained evaluation of policy performance and gain deeper insights into failure cases, we designed a task-specific evaluation rubric. Table XIV displays our rubric that the evaluator filled out when rolling out different policies. Take the DP policy as an example, the results in Table XIV demonstrate that H2R-augmented visual representation models not only improve overall success rates in real-world tasks, but also allow to accomplish more than half of the task consistently.
Appendix C Additional Ablation Study in Simulation Experiment
@@
From Table XIII, we observe that the encoder trained on H2R processed data shows consistent improvements across various tasks compared to the encoder trained on the original data, with the average success rate on all tasks ranging from 0.9% to 10.2%. Especially for the more challenging MoveCan task, it can improve the success rate by 25.5%. Additionally, while encoders trained on the relatively simple CutMix data show improvement on tasks in Robomimic, their performance in the PushT task remains slightly worse than the encoders trained on original data. These results demonstrate the effectiveness of using the robotic arm to cover the human hand in video data, as well as the effectiveness of H2R in imitation learning.